SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

DCZ CONFIGURATION STATUS

 
Run Number (1st Run After Change)
 
Date
 
Change
run 77299+ 14 Dec 07 fang - installed libL1DctApp.o in release 17.2.0.
run 69443+ 11 Jan 07 fang - installed libL1DctApp.o in release 16.0.0.
run 67475+ 9 Aug 06 fang - installed libL1DctApp.o in release 13.11.1.
run 66999 25 Jul 06 fang - validated libL1DctApp.o in release 13.11.1 with the cosmics.
run 65770+ 16 Jun 06 fang - linked back libL1DctCalApp.o.V01-01-06_on13.9.0.L1DctOdf.V01-01-10.
run 65769 16 Jun 06 fang - validated the ROM library libL1DctApp.o.13.11.0 in release 13.11.0.
runs 64258+ 27 Apr 06 fang - changed the ROM library to libL1DctCalApp.o.V01-01-06_on13.9.0.L1DctOdf.V01-01-10. This has the dependency on L1DctCalOdf and L1DctTeststand. The tags are: L1DctOdf V01-01-10, L1DctOnline in the release, L1DctCalOdf jtb_3nov04, L1DctCalOnline jtb_9mar05, L1DctDiagOnline jtb_1nov04, L1DctTeststand jtb_8feb05.
runs 64201+ 25 Apr 06 fang - changed the ROM library to libL1DctApp.o.V01-01-10_on13.9.0. It is built with online release 13.9.0. L1DctOdf V01-01-10 has taskDelay(2). The TSF firmware used, test_9 version, doesn't need taskDelay(). This library doesn't dependent on L1DctCalOdf and L1DctTeststand. The starup script is symbolically linked to startup-L1DctApp.ffang. The old startup was linked to startup-september_jamie. The same is done for the old DCT, although is not running, and name of the old startup is changed to startup-L1DctCalApp.
runs 64071, 64072, 64074, 64075 21 Apr 06 fang - these cosmic runs are used to test libL1DctApp.o.ff_13Apr06_on13.9.0. This library doesn't dependent on L1DctCalOdf and L1DctTeststand. Switched back to the original library when the test was done.
runs 63833+ 13 Apr 06 fang - changed the ROM library to libL1DctCalApp.o.ff_13Apr06_on13.9.0. This tag has taskDelay(60), which is 1 sec.
runs 63774+ 11 Apr 06 fang - changed the ROM library to libL1DctCalApp.o.ff_11Apr06_on13.9.0. This tag has taskDelay(30), which is 0.5 sec. The reason for adding extra delay is
runs 63349+ 27 Mar 06 fang - libL1DctCalApp.o.V01-01-09_on13.9.0 (release 13.9.0, L1DctOdf V01-01-09). taskDelay(10).
runs 62215+ 13 Feb 06 fang - changed the ROM lib to libL1DctCalApp.o.V01-01-07.on13.9.0 (release 13.9.0, L1DctOdf V01-01-07). Changed the out of frame print out message to include the element id.
run 61622+ 24 Jan 06 fang - changed the ROM lib to libL1DctCalApp.o.V01-01-06_on13.9.0
run 61621 24 Jan 06 fang - tested libL1DctApp.o.V01-01-08_on13.8.1_test built against release 13.8.1 and L1DctOdf V01-01-08.
runs 61585+ 22 Jan 06 fang - switched back to libL1DctCalApp.o.V01-01-06_on13.8.1
runs 61555+ 21 Jan 06 fang - changed the lib to libL1DctCalApp.o.V01-01-06_on13.9.0 built against online release 13.9.0. The tags are: L1DctCalOdf jtb_3nov04, L1DctCalOnline jtb_9mar05, L1DctDiagOnline jtb_1nov04, L1DctOdf V01-01-06, L1DctOnline V00-13-00, L1DctTeststand jtb_8feb05
runs 60913, 60916 4 Jan 06 fang - tested again libL1DctApp.o_on13.8.1_test. Run 60916 looked fine. A dch/dct calibration was taken and looked fine. Switched back to the original lib when the test was done.
runs 60465 - 60466 21 Dec 05 fang - tested libL1DctApp.o_on13.8.1_test based on release 13.8.1. The dependency on test stand lib is removed. Switched back to the original lib when the test was done.
runs 59377+ 14 Nov 05 fang - changed the library to libL1DctCalApp.o.V01-01-06_on13.8.1, built against online release 13.8.1, and there is NO change in the tags from release 13.3.0.
runs 58962 - 58974 4 Nov 05 fang - libL1DctCalApp.o_13.8.0_test, libL1DctCalApp.o_13.8.0_verbose_test are used for these cosmic runs to test a new version of TSF firmware. In this version all unphysical segment data are save by L1DctNewTsfL1Accept.cc to superlayer=10, celle=22 (libL1DctCalApp.o_13.8.0_test), and data are printed out (libL1DctCalApp.o_13.8.0_verbose). Switched back to the original lib when the test was done.
runs 58928+ 3 Nov 05 fang - changed the library to libL1DctCalApp.o.V01-01-06_on13.8.0, built against online release 13.8.0, and there is NO change in the tags from release 13.3.0.
runs 57205+ 16 Aug 05 fang - changed the library to libL1DctCalApp.o.V01-01-06_on13.4.0, built against online release 13.4.0, and there is NO change in the tags from release 13.3.0.
runs 56740+ 2 Aug 05 fang - changed the library to libL1DctCalApp.o.V01-01-06_on13.3.0, built against online release 13.3.0 using these tags: L1DctCalOdf jtb_3nov04, L1DctCalOnline jtb_9mar05, L1DctDiagOnline jtb_1nov04, L1DctOdf V01-01-06, L1DctOnline V00-12-01, L1DctTeststand jtb_8feb05
runs 50570-50578 29 July 04 everyone - switched tsf firmware to 2_12. so are gonna run with glt fed from zpd and old blt system. putting in new glti. new l3 build with trg config etc.. (we are gonna run glt from new system but l3 from old tsf segments so all dcz data is still going to be in 'parasite' tc's). first runs are l1passthru some changing of the trg configuration...
runs 50538+ 28 July 04 jamie - ok trying marcs latest attempt (blt_testing2/blt_0) firmware on the tsfs. this has fixed the lack of driving the zpd. but looking at the data (after) shows that some tsf daq data (& zpd & blt paths) is missing for certain parts of y boards.
runs 50526+ 28 July 04 jamie - hmmm. look like the new daq offset is right. but the blt doesnt seem to be driven properly with this new firmware. im now trying all the new sets at the same time (blttests/blt_0,1,2,3 in 4 different quadrents of the detector in phi).
runs 50506+ 27 July 04 jamie - changed the debug blt daq lat from 1998 to 1997 to see if this aligns the blt debug and default data in time after the tsf firmware changes.
runs 50504+ 27 July 04 jamie - changed the tsf firmware to blttest/blt_0 which should have the same blt output phasing as 2_10 and also the tsf daq improvements of 2_12. this seems to work for the blt and tsf daq but doesnt seem to feed the zpd correctly - we leave it in anyway to check it works for the blt as planned.
runs 50480+ 26 July 04 jamie + su dong - changed the tsf firmware to 2_10. this seems to have the right timing on the blt output to drive the new blt. but it has bad timing for tsf daq data - so this looks screwed up.
runs 50468+ 26 July 04 jamie + su dong - put the BLT into the system (slto 2 of ZPD crate). moved ZPDs around to allow this to happen. ZPD crate now looks like:
 
middle - zpd
SLOT   0    1    2    3    4    5    6    7    8    9    10    11    12    13    14    15
SLACID       (BLT)2211          4508 4509 4511      4507 4505 4501        4506        4502
SECTOR           -              0    1    2         3    4    5           6           7
INTNUM         3506             70   80   76        78   75   74          72          73
runs 50384+ 23 July 04 jamie - switched the tsf in slot 0 (X0) which was prod board 10 for prototype with red dots (unnumbered). this is because it looks like this board has one zpd output channel hot. (seems to have fixed the problem).
runs 50359-50360 22 July 04 jamie/eunil/su dong/xuedong - recabled bad tsf->zpd cables (a number). replaced 2 zpdi's (slot 1 and 3) this gave data to these boards. switched tsfi 51 -> 66. switched zpd boards in slot 1 (80->69) and slot 3 (78 -> 71). switched zpd boards in slot 6 and slot 7.(actually this was after the above runs).
runs 50344-50349 22 July 04 jamie/eunil/su dong/xuedong - put in new crate. put in 24 tsf boards and cabled up tsfi's & zpdi's. crates look like: top - tsf y
 
SLOT   0    1    2    3    4    5    6    7    8    9    10    11    12    13    14    15
BOARD       26        25        21        8              27    29          P4    P1
SLACID      4536      4535      4531      4518           4537  3554        -     -
SECTOR      0         7         1         6              2     5           3     4
INTNUM      60        64        59        61             66    62          52    58

middle - zpd
SLOT   0    1    2    3    4    5    6    7    8    9    10    11    12    13    14    15
SLACID      4509      4508      4511      4507      4505       4501        4502        4506
SECTOR      0         1         2         3         4          5           6           7
INTNUM      70        80        76        78        75         74          72          73

bttm - tsf x
SLOT   0    1    2    3    4    5    6    7    8    9    10    11    12    13    14    15
BOARD  10   23   5    17   7    28   4    20   6    P3   13    11    15    30    18    14
SLACID 4520 4533 4515 4527 4517 3553 4514 4530 4516 -    4523  4521  4525  3555  4528  4524
SECTOR 0    15   1    14   2    13   3    12   4    11   5     10    6     9     7     8
INTNUM 45   65   53   63   46   43   48   42   49   38   50    37    54    56    57    55
5:20pm 19 July 04 jamie - so we ran all weekend with this daqtest/test8 firmware and the data looks ok. i have now removed all the tsfs to get the handles put on them. eunil has removed the zpdis to get the clamps put on them. hopefulyl we will reapper in ir2 with 24 (tested) tsfs and 2 crates in the next couple of days.
all day! 16 July 04 jamie - 10 new sets of firmware to test from marc. test8 seems to be the most successful so am leaving this in. (data looked good from a dch-dct calib with this firmware). so am taking colliding beam run (50157) with this firmware and if its ok maybe will take more data over the weekend with this.
5:30pm 14 July 04 jamie - trying to test marcs daq_test firmware versions (1,2,3,4). test 1. have left 2 in the system and may test with a parasitic run. will try 3,4 tomorrow. tested others next day but all were - unsuccessful.
3pm 14 July 04 jamie - switched boards in slot 1 and slot 3 with each other. also switched boards in slot 0 and slot 9. (this is because was getting data corruption reported in rom logfile for slots 1 and 9 want to see if this moves with the board). current slot / board numbers are: 0/23/x9, 1/29/y0, 2/p1/y1, 3/27/y2, 4/p4/y3, 5/prd/y4, 6/11/x0, 7/8/x1, 8/25/x2, 9/30/x3, 10/21/x4, 11/26/x5, 12/15/x6, 13/10/x7, 14/p3/x8
50090 14 July 04 jamie - changed TSF firmware back to version 2_6 (from 2_10) to see if this effects the tsf daq data corruption
50063 13 July 04 jamie - switched boards in slot 11 and slot 9 eunil - fixed x4 miscabling of zpd3 (tsfi side). su dong - replaced tsfi x4 with production panel. turned the latches around for some of the tsfis
6 pm + 9 July 04 jamie/eunil/xuedong - added 15 tsfs, 15 tsfi's, 8 zpd, 8 zpdis. tsf boards are as follows slot/board no./ board. 0/30/x9, 1/27/y0, 2/p1/y1, 3/29/y2, 4/p4/y3, 5/prd/y4, 6/11/x0, 7/8/x1, 8/25/x2, 9/26/x3, 10/21/x4, 11/23/x5, 12/15/x6, 13/10/x7, 14/p3/x8
- 7 July 04 Eunil and Xuedong: Found again (!) a zpdi cable (blue) was stepped below the latch of the zpdi (The label was TSF-Y01-2). Not sure the cable was damaged or not.
1:20am+ (next reconfigure) 7 July 04 jamie - put in 6 new production tsfs which just arrived. tested setup with a parasitic run which worked ok. the boards are in the following way: slot 1 #27 (y1), slot 3 #29 (y2) slot 6 #30 (x2) slot 9 #25 (x3) slot 11 #26 (x3) slot 13 #23 (x4).
49699 2 July 04 jamie - changed zpd track prescale factor to 1. this fixed the apparent drop in zpd efficiency (to ~0.1).
5:00pm+ (49672) 1 July 04 jamie & eunil - redid neighbour cabling for TSFi X4 (was wrong again). changed TSF firmware to 2_6 (a backwards step) to try to understand ZPD efficiency drop.
5:30pm+ (49630) 30 June 04 jamie - moved system to ZPD (from BLT). switched TSFi in X4 and X5 (to try to trace transmission problem in X4). redid neighbour data cabling for X2/X3/X4. TSF firmware still blt_3 (seems to have right phasing for BLT). wedge 1,2 still covered by system.
10:22am+ (49618) 30 June 04 jamie - changed TSF firmware to blt_3 (change in blt output phasing).
1:50pm+ (49593+) 29 June 04 jamie - changed TSF firmware to blt_2 (change in blt output phasing) will also try ( blt_3 tomorrow).
6pm+ (49568+) 28 June 04 jamie - put BLT in (instead of ZPDs) and changed TSF firmware to blt_1 (change in blt output phasing) will also try (blt_2 / blt_3 tomorrow).
2:12pm+ (49465+) 25 June 04 so TSF board X5 seems to be flaky sometimes (Sl8,1,2 have been sometimes bad soemtimes good). eunil - jamie, replaced BLT with ZPDS covering wedges 1,2.
10:39am+ 24 June 04 switched TSF boards X4 and X5 to see if bad X4 SL8 data moves. (board X4 was red dots, X5 was #3 - from now on swapped). tried to see if this swap fixed/moved TSF X4 problem with DCH-DCT calib. but the problem didnt show up on the calibration before the switch (calib 3726 or after calib 3727).
49405 23 June 04 updated TSF firmware to 2_10 version. this has different BLT timing and also fixes MC with TSFi problem. - data after this shows a problem with TSF X4 SL 8 (see page 9 /BFROOT/www/Detector/Trigger/upgrade/fastmon//boyd-DczFastMon-0049418-20040624-044600.ps.gz ).
2:15pm 17 June 04 switched wedge to cover sectors 1, 2 and took out ZPDs and put in BLT. tested different TSF firmware (changing the clk30 that the tsf->blt data is shipped out on). will now try to take cosmic run with blt in the system and with 2_6 firmware. - jamie
49188 16 June 04 switched wedge to cover sectors 3, 4 (so Zpds 2,3,4,5 are being fed with tracking possible in 3 & 4). the purpose is to scan for more problem areas in phi (like X14). zpd mask maybe wrong.
49164 15 June 04 switched wedge to cover sectors 5, 6 (so Zpds 4,5,6,7 are being fed with tracking possible in 5 & 6). the purpose is to scan for more problem areas in phi (like X14). zpd mask maybe wrong.
49135 14 June 04 problem with new firmware giving corrupted data for tsf (get lots of fex errors). switch back to 2_6 firmware.
49131 14 June 04 Jamie: Switched the TSFi's in slots 6 and 9 (corresponding to X14 and X15). Also updated TSF firmware to 2_8 which has a fix to the MC problem (when TSFi is present). MC test confirmed firmware works ok and parasite run confirms it runs ok in data taking mode.
49054 12 June 04 Jamie: Switched New TSF board X14 and X15 (properly this time) also switched TSFi's X0 and X1. tested that all was ok with a parasitic run.
49006 10 June 04 jamie&eunil: switched TSFi 20004439 with 20004445 as it was not sending frame bits through to ZPD (although was receiving frame bits). was causing glink errors but now seem to be fixed.
49005 10 June 04 Jamie: Switched New TSF board X14 and X15 to see if X14 problem moves with board - actually stupidly swapped boards X0 and X1 instead
48983 9 June 04 Eunil/Jamie: Moved wedge to cover sectors 0,7
(1st run after rod on 03-Jun) 3 June 04 Eunil: Install a new zpd firmware has all machinery for the double buffering for daq memories but not actually in action.
48731 (1st reconfigure after) 2 June 04 jtb: last 4 tsf daq lat changes didnt do anything because i was changing tsfConfig.txt but actually need to make an xtc file (dctNewTsfConfig.xtc) from this text file. this explains why i was having problems centering the data. so i have now changed the xtc file to have param 36 need to wait for next configure for it to take effect though.
48725 (1st reconfig after) 2 June 04 jtb: changed tsf daq latenmcy to 33 (from 36) to try to get tsf tick distbns to agree between old and new tsfs (this is begining to be annoying!)
48707+ (1st reconfig after) 1 June 04 jtb: changed tsf daq latenmcy to 36 (from 34) to try to get tsf tick distbns to agree between old and new tsfs
48703+ 1 June 04 jtb & eunil: put 4 ZPDs in and took BLT out (zpd cabled to be consistent with TSFs). While we were doing this, we made a mistake of zpdi bottom lever arm stepping on a zpdi cable ("the blue cable" labeled as TXZ-03T3-Z2X1). Not clear there was a damage or not.
48547-48572 28 May 04 jtb: changed BLT latency for parasite BLT (in ZPD crate) had to hack the DCZ library for that to work (now DCZ and DCT boot with different libraries). increased lat to 1998 (+6) to center BLT data. (these are colliding beam data with parasite & production BLT in system).
48498-48506 26 May 04 jtb: put BLT into ZPD crate. cosmic runs. now 6 TSFs TSF X2,3,4,5 / Y1,2 (moved wedge).
48278 19 May 04 jtb: changed the tsf firmware to proto_2_6 (with improved neighbour data stuff).
> 48343 21 May 04 sjb: Upgraded ZPDs to v10. This includes nSegMiss cut and tweak to Fitter firmware to not include hitmap in tracks with failed fits.
??? 17 May 04 ew, sjb: Added 5 more ZPDs to make a full 8. Edited teststand_map/bltZpdMap.txt for new ZPDs, teststand_map/newTsfxMap.txt because 1 TSF is back at the teststand, and June2003/zpdLutList.txt to not read in the ZPDi mask for now. We will probably need to recheck all this config stuff before going back in the partition.
47853 (or 1st config after) 6 May 04 jtb: changed tsf daq lat to match old and new tsf tick (as much as possible) after new firmware reduced glink latency by 1 clk4
47595 28 Apr 04 jtb: Add DCZ in the partition to test new TSF firmware (version 2_5) which should reduce TSF glink latency by 1clk 4. also changed zpd config so that zpd data is centered (at tick=3) after this change (17 -> 18) also removed the data prescaling.
47562 26 Apr 04 ew: Add DCZ in the partition to test sjb_20Apr04. (All prescale=10, and input prescale=1)
unknown 20 Apr 04 sjb,ew: Updated libL1DctCalApp.o to fix various bugs introduced along with (all/input) prescaling (Tag: sjb_20Apr04 for L1DctOnline and L1DctOdf).
unknown 6 Apr 04 sjb: Updated ZPD firmware to version 9 for better UserReset stability.
unknown 6 Apr 04 jtb, ew: Re-installing new TSF boards to the crate, after a fix up of the power tray of the TSF crate (board 1 in slot 1 2 in 6, and 3 in 9). Power on new crate seems stable. A piggyback run showed no trouble on zpd distribution.
46444 24 March 04 jtb, sjb, ew: Lots of fun with master crate power cycles. Upgraded ZPD firmware to v07 which will hopefully fix it by resetting the FF DCMs upon UserReset.
46188 17 March 04 Eunil: Changed zpd firmware to version 6 to fix daqformat=2 trouble. Also set the prescale for input mask (zpd) to be 10 as a default value for now.
46013 11 March 04 Jamie: Changed tsf firmware to version 2_3 this (hopefully) stops the jumping around by a clk4 of the tsf data. also Eunil & i updated the zpd daq offset to move the peak back to 3 (changed both input an trk offsets to 17).
45656 4 March 04 Jamie&Marc: Changed tsf firmware to reduce the latency. This looked like it would take 2 clk4s off - but now looks like it may oscillate taking 1 or 2 depending on the glink frame bit posn. we will leave this in over night and look at the data after that to ssee whats going on. if we were to leave this (or any latency reduced) firmware in we will need to change the tsf & zpd daq offsets.
45613 3 March 04 Stephen: Upgraded ZPD firmware to v5; upgraded L1DczCalApp to tag sjb3Mar04 which has improved ZPD consistency checks after configuration.
45527-45535 1 March 04 Changed the DCZ online library to one in which the zpd input data is prescaled by 10. this is experimental and the DCZ offline code may not be able to run on the xtc file produced - we will see?
45387-8 27 Feb 04 Changed the LUT loaded into the NEW TSFs so that it has a reduced delay (by 2clk 8s mostly) this is just a test to see what effect this has on ZPD eff. (the LUT loaded is in /dataflow/rdf/Dct/June2003/tsfLutTC.xtc_reduced_delay). jamie
45389+ 27 Feb 04 Changed the LUT loaded into the NEW TSFs back to the normal one. (the LUT loaded is in /dataflow/rdf/Dct/June2003/tsfLutTC.xtc_original). jamie
45387-8 27 Feb 04 Changed the LUT loaded into the NEW TSFs so that it has a reduced delay (by 2clk 8s mostly) this is just a test to see what effect this has on ZPD eff. (the LUT loaded is in /dataflow/rdf/Dct/June2003/tsfLutTC.xtc_reduced_delay). jamie
45279 24 Feb 04 Stephen: Updated ZPD firmware to v4 which should be more robust against loosing the clock. Updated L1DctCalApp.o_sjb to include UserReset before autodetection.
44822 13 Feb 04 Changed TSF DAQ Latency back from 34 to 35. as this is the best value for old tsf new tsf daq data agreement.
44743 11 Feb 04 put TSF boards back. order is slot 0 Y0 #1, slot 6 X0 #2, slot 9 X1 preseries bd. also updated firmware on tsf's to proto_2_2 for #1, #2 and production_2 for preseries board.
44678 9 Feb 04 switched tsf boards 1 (Y0 -> X1) and 2 (X1 -> Y0) so now they are slot 0 Y0 #2, slot 6 X0 #4, slot 9 X1 #1.
first physics after 44527 5 Feb 04 Changed TSF DAQ Latency from 35 to 34. this is to see if the tsf daq data tick now oscillates around as the zpd input data does. (jamie).
44328+ 29 Jan 04 Switched out the preseries board for neighbour data tests - put in tsf # 2 instead in slot 9 (X board) now the crate looks like slot 1 tsf #1 Y, slot 6 tsf # 4 x, slot 9 tsf #2 X.
44325->26 29 Jan 04 Switched the position of the new preseries TSF board so its in slot 9 (X board) now the crate looks like slot 1 tsf #1 Y, slot 6 tsf # 4 x, slot 9 newpreseries board X.
44311->44317 29 Jan 04 added eunil's new firmware to zpd's (also changed zpd config daq offset for input and tick data from 18->16 all in decimal) see HN 554 for more details. Also switched the position of the new preseries TSF board so its in slot 6 (X board) now the crate looks like slot 1 tsf #1 Y, slot 6 newpreseries board X, slot 9 tsf #4 X.
44277 -> 44291 27 Jan 04 put in the new preseries tsf board in slot 1 (as tsf y) so arrangment of tsf was Y (slot 1) preseries, X (slot 6) tsf #2, x (slot 9) tsf #4.
44205 26 Jan 04 Switched TSF X board sin slots 6 & 9 somehow this cured a MC segment dropping problem with board 2 in slot 6(?) se hn 551.
43993 19 Jan 04 DCZ IN AND WORKING
43722-23 12 Jan 04 DCZ IN but maybe problems??
43598-601 9 Jan 04 DCZ IN - but many problems related to master crate reboot we think
43529-40 8 Jan 04 DCZ IN
43179 15 Dec 03 thats the end of DCZ running for 2003 (for various annoying reasons - dont ask)
43104 12 Dec 03 put the preseries tsf board in as tsf x2 (slot 6) (jamie)
42777 3 Dec 03 removed ip correction for zpd. (jamie, eunil & xuedong)
42359 21 Nov 03 replaced all 3 TSFis with new production set. see HN 530 for details (jamie & xuedong)
42275 18th Nov 03 Loaded into ZPDs IP correction LUT. see HN 524 for details (stephen)
42263 17th Nov 03 Changed online DCZ library to one based on 10.10.0. see HN 523 (jamie & gerald)
42120 (fixed zpdiMask after run 42141
(i think - in HN it says def fixed by 42265))
12th Nov 03
(fixed zpdi mask after 13 Nov 03)
changed tsf wedge to sector 0,1 (from sectors 2,3). see HN 520. (jamie)
(forgot to update zpdiMask correctly this was fixed by stephen & kris later in the meantime zpd data was crap.).



PLEASE ADD CHANGES SUCH AS

Changes to physical hardware (ie. new/different boards)
Changes to firmware
Changes to configuration params (inc. LUTs / DAQ offsets etc..)
Changes to online software
Reference details in HN where the number given is relative to:
http://babar-hn.slac.stanford.edu:5090/HyperNews/get/trignews/

OTHER IR2 DCZ ACTIVITY

Please log other activity where you temporarily changed the DCZ configuration and then (hopefully) restored it to its previous state.
Date & Time Who Activity
5 Feb 04;
14:30 - 16:30
Stephen Tried beta version of ZPD firmware v4. It didn't work, so go back to v3.
6 Feb 04;
10:30 - 12:30
Stephen More firmware testing; swapping ZPDs around. Left in orig state.